Pls write python code

1. We want to obtain the volume contained between a sphere of radius 0.75 and another larger sphere of radius1. Modify the Monte Carlo written for the volume of a sphere. Use N=10,000 sampling points (random points in the cube) ax
Answer #1

Below is the python code to calculate the volume between two spheres:

r1= 0.75, r2= 1

pi = 3.14

Vol1= 4.0/3.0*pi* r1**3

Vol2= 4.0/3.0*pi* r2**3

v= Vol2 - Vol1

print('The volume between the spheres is: ',v)